in a region uniform electric field E=10^3 N/C exists along +y axis .a charged particle (mass =1g, charge =1mC) enters with velocity 10 m/s along +x axis find speed of the particle when displacement along x axis is 10 cm.

Grade:12th pass

2 Answers

Shailendra Kumar Sharma
6 years ago
Assume the particle is moving along X-axis and Along +Y Axis there is an Electric field. It can be assumed a gravitational field. 
so the force on object will be qE and acceleration along +Y axis will be qE/m 
a=qE/m
a= {1x10-3 * (103)}/(10-3) =103 m/s
Now velocity when displacement along x-axis 10cm. 
time taken in 10 cm travel will be 10cm/10m/s = 0.01s. 
So the y component of velocity will be Vy=at =1000*(.01) =10m/s
So the velocity of particle will be \sqrt{100+100} =10\sqrt{2}  or 14.14 m/s
ankit singh
3 years ago
in a region uniform electric field E=10^3 N/C exists along +y axis .a charged particle (mass =1g, charge =1mC) enters with velocity 10 m/s along +x axis find speed of the particle when displacement along x axis is 10 cm.